### v1

# Lieb-Schultz-Mattis-Type and Laughlin-Type Argument for the Quantum Hall Effect in Lattice Fermions with Spiral Boundary Conditions

Abstract
We derive the condition for the occurrence of the quantum Hall effect in two-dimensional lattice systems, expressed as $\phi\nu-\rho\in\mathbb{Z}$, where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively. By employing spiral boundary conditions, which treat the system as an extended one-dimensional chain, this condition is obtained directly through a Lieb-Schultz-Mattis-type and Laughlin-type argument. This approach improves upon the preceding work based on conventional periodic boundary conditions, where the condition was derived indirectly containing redundant system-size dependence. The key to this approach is that the spatial directions of the external force and the response can be manipulated by a factor of the system size.
–Introduction–
The quantization of the Hall conductance in two-dimensional (2D) electron systems, known as the quantum Hall effect (QHE), is one of the most remarkable manifestations of topology in condensed matter physics. The quantized Hall conductivity is expressed in terms of the Chern number of the occupied bands[1, 2, 3]. From the viewpoint of many-body theory, the quantization of the Hall conductance can also be derived from Laughlin’s gauge argument,[4] which explains the integer change of polarization upon adiabatic flux insertion. In lattice systems, the condition for the appearance of the QHE is given by a Diophantine equation that relates the magnetic flux, the Chern number, and the electron density.[1, 5]
On the other hand, a similar topological constraint was discussed by Lieb, Schultz, and Mattis (LSM) for one-dimensional (1D) quantum spin systems,[6] and later generalized to various correlated systems[7, 8, 9, 10] and higher dimensions[11, 12, 13, 14, 15, 16]. These approaches are closely related through the concept of the modern theory of polarization introduced by Resta[17, 18, 19]. The theory of polarization has also been applied to various systems[20, 21], and also to the calculation of topological numbers of 2D lattice systems by applying Laughlin’s arguments of flux insertion[22].
Lu, Ran and Oshikawa derived the conditions for the appearance of the QHE in lattice systems with interactions by combining the LSM-type argument with Laughlin-type argument as[23, 24]
$\phi\nu-\rho\in\mathbb{Z},$ (1)
where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively. In their theory based on periodic boundary conditions (PBCs), the condition for QHE appears indirectly passing through the intermediate relation which depends on an artificial geometric parameter. Therefore, improvement of this proof is desirable to eliminate this redundancy.
In this Letter, we reformulate the argument using spiral boundary conditions (SBCs)[16, 25, 26, 27], which allow a 2D system to be regarded as an extended 1D chain (see Fig. 1). By employing the LSM-type and Laughlin-type arguments under the SBC, we derive the condition for QHE (1) directly, without invoking any system-size dependence. This formalism provides a unified and more transparent understanding of the QHE from the standpoint of symmetry and topology.
In the rest of this Letter, we explain the argument by Ref. [23]. Then we reformulate this argument based on SBCs and derive Eq. (1). Throughout this letter, the electron charge, the reduced Planck constant, and the lattice constant are set to be unity as $e=\hbar=a=1$.
Figure 1: Several boundary conditions for 2D square lattices with $L_{1}\times L_{2}$ sites: (a) Conventional periodic boundary conditions (PBCs), (b) Spiral boundary conditions (SBCs). For the systems with SBCs, the lattices are labeled as extended 1D chains. $\bm{A}_{i}$ and $\bm{A}_{i}^{\prime}$ are primitive vectors.
–Conventional periodic boundary conditions–
First, we discuss how the condition given by Eq. (1) is derived following the discussion by Lu, Ran and Oshikawa[23]. The Hamiltonian for the fermions in the 2D square lattice with $L_{1}\times L_{2}$ sites under periodic boundary conditions is given by[28]
$\mathcal{H}=-t\sum_{m,n}(c^{\dagger}_{m+1,n}{\rm e}^{{\rm i}\theta_{m,n}^{1}}c^{\mathstrut}_{m,n}+c^{\dagger}_{m,n+1}{\rm e}^{{\rm i}\theta_{m,n}^{2}}c^{\mathstrut}_{m,n}+\mbox{H.c.})+V(\{n_{m,n}\}),$ (2)
where $c^{\mathstrut}_{m,n}$ ($c_{m,n}^{{\dagger}}$) is the annihilation (creation) operator at $(m,n)$ site, $t$ is the hopping amplitude, $V(\{n_{m,n}\})$ is the fermion interactions consists of the number operator $n_{m,n}\equiv c_{m,n}^{{\dagger}}c_{m,n}^{\mathstrut}$. The gauge field is introduced in the Landau gauge as
$\theta^{1}_{m,n}=0,\qquad\theta^{2}_{m,n}=2\pi\phi m,$ (3)
which give the magnetic flux $B=-\frac{2\pi\hbar}{ea^{2}}\phi$ with
$\phi=\frac{1}{2\pi}\sum_{\Box}\theta_{m,n}=\frac{p}{q},$ (4)
where $p$ and $q$ are coprime integers. The summation for the plaquette is taken as shown in Fig. 2(a),
$\sum_{\Box}\theta_{m,n}=\theta_{m,n}^{1}+\theta_{m+1,n}^{2}-\theta_{m,n+1}^{1}-\theta_{m,n}^{2}.$ (5)
Figure 2: The gauge field $\theta^{\alpha}$ in a square lattice in a magnetic field with (a) periodic boundary conditions, and (b) spiral boundary conditions.
In the absence of a magnetic field ($\phi=0$), the Hamiltonian possesses translational symmetry by one lattice spacing along both directions, $\mathcal{H}=T_{i}^{\mathstrut}\mathcal{H}T_{i}^{-1}$. For $\phi\neq 0$, this translational symmetry is broken due to the position dependence of $\theta^{2}_{m,n}$ and sublattice structure appears as shown in Fig. 3. Therefore, $L_{1}$ is selected so that $L_{1}/q\in\mathbb{Z}$. On the other hand, a similar symmetry is recovered as $\mathcal{H}=\tilde{T}_{i}^{\mathstrut}\mathcal{H}\tilde{T}_{i}^{-1}$ by introducing the magnetic translation operators
$\tilde{T}_{1}=U_{2}^{L_{2}\phi}T_{1},\qquad\tilde{T}_{2}=T_{2}.$ (6)
The polarization operator
$U_{i}=\exp\left[{\rm i}\frac{2\pi}{L_{i}}\sum_{\bm{x}}x_{i}n_{\bm{x}}\right].$ (7)
with $\bm{x}=(x_{1},x_{2})$ satisfies the relation
$T_{1}U_{1}T_{1}^{-1}={\rm e}^{-{\rm i}2\pi L_{2}\hat{\rho}}U_{1}.$ (8)
where $\hat{\rho}$ is the density operator, and $L_{2}$ represents the cross section of the 2D system. As in the LSM argument for $d\geq 2$, $L_{2}$ and $q$ should be chosen to be coprime.
Figure 3: Sublattice structure of the system $(L_{1},L_{2},q)=(6,4,3)$ with a magnetic field ($\phi=p/q$). Here, $L_{1}/q$ is chosen to be an integer.
Now we consider the quantum Hall effect in this lattice system. The quantized Hall conductivity is characterized by the Chern number $\nu$ and the density $\rho$ as
$\sigma_{xy}=\frac{e^{2}}{h}\nu,\qquad\rho=\frac{r}{q},$ (9)
where $r$ is an integer. Since the magnetic flux $\phi=p/q$ induces a sublattice structure with $q$ sites as shown in Fig. 3, a commensurate filling is required to open an energy gap. Thus, for the QHE where $\nu$ is an integer, the following condition must hold:
$\phi\nu-\rho=n,\qquad n\in\mathbb{Z},$ (10)
which is equivalent to the Diophantine equation
$r=\nu p+nq.$ (11)
[…]
$z(\Phi)=z(0){\rm e}^{-{\rm i}\nu\Phi},\quad z(\Phi_{2})=\braket{\Psi_{2}(\Phi_{2})|U_{1}|\Psi_{2}(\Phi_{2})},$ (12)
leading to
$\nu=-\frac{1}{\Phi}\mathrm{Im\,ln\,}z(\Phi).$ (13)
The adiabatic flux insertion is expressed as
[…]
with
$F_{i}(\Phi_{i})=\mathcal{T}\exp\left[-{\rm i}\int_{0}^{T}\mathcal{H}_{i}(\Phi_{i}\tfrac{t}{T}){\rm d}t\right],$ (15)
where $\mathcal{T}$ denotes time ordering, $T$ is a sufficiently long period, and $\mathcal{H}_{i}(\Phi)\equiv U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$.
Equation (12) then yields the operator relation
$F_{2}(\Phi)]^{-1}U_{1}F_{2}(\Phi)=U_{1}{\rm e}^{-{\rm i}\nu\Phi}.$ (16)
This is also an expression of the Laughlin’s argument for QHE. Since $F_{i}(2\pi)$ plays a similar role to $U_{i}$, we can introduce a magnetic translation operator as an analogy of Eq. (6),
$\bar{T}_{1}=F_{2}(2\pi L_{2}\phi)T_{1},$ (17)
which satisfies $\mathcal{H}=\bar{T}_{i}^{\mathstrut}\mathcal{H}\bar{T}_{i}^{-1}$ for $\phi\neq 0$. We then obtain
$\bar{T}_{1}U_{1}\bar{T}_{1}^{-1}=U_{1}{\rm e}^{{\rm i}2\pi L_{2}(\phi\nu-\rho)}.$ (18)
Thus, following an LSM-type argument, the condition for $z(\Phi)\neq 0$ becomes
$z(\Phi)$ $=\braket{\Psi_{2}(\Phi_{2})|\bar{T}_{1}U_{1}\bar{T}_{1}^{-1}|\Psi_{2}(\Phi_{2})}$ (19)
$=z(\Phi){\rm e}^{{\rm i}2\pi L_{2}(\phi\nu-\rho)}.$ (20)
Hence,
$L_{2}(\phi\nu-\rho)\in\mathbb{Z},$ (21)
Since this must hold for arbitrary integer $L_{2}$, we finally obtain the condition $\phi\nu-\rho\in\mathbb{Z}$ which is precisely the condition for the QHE given by Eq. (1). However, the appearance of the artificial parameter $L_{2}$ is undesirable, because $L_{2}$ must be chosen to be relatively prime to $q$ so that a redundancy remains.
Figure 4: Fluxes $\Phi_{i}$ for 2D lattice applied as $\mathcal{H}_{i}(\Phi)=U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$. In the system with SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect of the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation $U_{1}^{\prime}=U_{2}^{\prime L_{2}}$.
–Spiral boundary conditions–
Next, we reconsider the above discussion under spiral boundary conditions (SBCs), where the Hamiltonian (2) is rewritten as a 1D model with long-range hopping,
$\mathcal{H}^{\prime}=-t\sum_{j=1}^{L}(c^{\dagger}_{j+1}{\rm e}^{{\rm i}\theta_{j}^{1}}c^{\mathstrut}_{j}+c^{\dagger}_{j+L_{1}}{\rm e}^{{\rm i}\theta_{j}^{2}}c^{\mathstrut}_{j}+\mbox{H.c.})+V(\{n_{j}\}),$ (22)
where $c^{\mathstrut}_{j}$ ($c_{j}^{{\dagger}}$) is the annihilation (creation) operator at $j$ th site of the extended 1D chain of SBCs, and $L\equiv L_{1}L_{2}$ is the total number of lattice sites. The gauge field is given by (see Fig. 2(b))
$\theta^{1}_{j}=0,\qquad\theta^{2}_{j}=2\pi\phi j.$ (23)
Now we introduce the polarization operators in SBCs. For this purpose, we reinterpret the polarization operators in PBCs (7). The primitive vectors in the real space (see Fig. 1(a)) and the reciprocal lattice space of 2D systems as
$\bm{A}_{1}$ $=(L_{1},0),$ $\qquad\bm{B}_{1}$ $=2\pi(\tfrac{1}{L_{1}},0),$ (24a)
$\bm{A}_{2}$ $=(0,L_{2}),$ $\bm{B}_{2}$ $=2\pi(0,\tfrac{1}{L_{2}}).$ (24b)
Then these vectors satisfy the following relation,
$\bm{A}_{i}\cdot\bm{B}_{j}=2\pi\delta_{ij}.$ (25)
The polarization operators (7) are written using the above primitive vectors in the reciprocal lattice space as
$U_{i}=\exp\left[{\rm i}\sum_{\bm{x}}\bm{B}_{i}\cdot\bm{x}n_{\bm{x}}\right].$ (26)
For SBCs as shown in Fig. 1(b), the primitive vectors are chosen so that they satisfy the relation (25),
$\bm{A}^{\prime}_{1}$ $=(L_{1},-1),$ $\qquad\bm{B}^{\prime}_{1}$ $=2\pi(\tfrac{1}{L_{1}},0),$ (27a)
$\bm{A}^{\prime}_{2}$ $=(0,L_{2}),$ $\bm{B}^{\prime}_{2}$ $=2\pi(\tfrac{1}{L_{1}L_{2}},\tfrac{1}{L_{2}}).$ (27b)
The polarization operators in 2D systems with SBCs are introduced as
$U^{\prime}_{i}=\exp\left[{\rm i}\sum_{\bm{x}}\bm{B}^{\prime}_{i}\cdot\bm{x}n_{\bm{x}}\right].$ (28)
For $i=1$, due to the relation $\bm{B}_{1}^{\prime}=\bm{B}_{1}$, $U_{1}^{\prime}$ is the same operator as $U_{1}$. The representation in SBCs is given by the relation
$\frac{2\pi jn_{j}}{L_{1}}=\frac{2\pi x_{1}n_{\bm{x}}}{L_{1}}\quad(\mathrm{mod}\,2\pi).$ (29)
For $i=2$, the exponent part of the polarization operator is confirmed as that of SBCs as
$\sum_{\bm{x}}\bm{B}^{\prime}_{2}\cdot\bm{x}n_{\bm{x}}$ $=\sum_{m=1}^{L_{1}}\sum_{n=0}^{L_{2}-1}2\pi\frac{m+nL_{1}}{L_{1}L_{2}}n_{m,n}$ (30)
$=\sum_{j=1}^{L}\frac{2\pi}{L}jn_{j},$ (31)
where $j=m+nL_{1}$. Thus the polarization operators for SBCs are identified as
$U_{1}^{\prime}=\exp\left[{\rm i}\sum_{j=1}^{L}\frac{2\pi jn_{j}}{L_{1}}\right],\quad U_{2}^{\prime}=\exp\left[{\rm i}\sum_{j=1}^{L}\frac{2\pi jn_{j}}{L}\right].$ (32)
In SBCs, the translation and polarization operators for different directions are related as
$T_{2}^{\prime}=T_{1}^{\prime L_{1}},\qquad U_{1}^{\prime}=U_{2}^{\prime L_{2}}.$ (33)
Then, the counterpart of Eq.(8) becomes
$T_{1}^{\prime}U_{2}^{\prime}T_{1}^{\prime-1}=T_{1}^{\prime}U_{1}^{\prime 1/L_{2}}T_{1}^{\prime-1}={\rm e}^{-{\rm i}2\pi\hat{\rho}}U_{2}^{\prime}.$ (34)
We choose $U_{2}^{\prime}$ rather than $U_{1}^{\prime}$ here because it eliminates the $L_{2}$-dependence from this relation. This was introduced for an extention of the LSM theorem to higher dimensions[16]. The magnetic translation operators are
$\tilde{T}_{1}^{\prime}=U_{2}^{\prime L_{2}\phi}T_{1}^{\prime}=U_{1}^{\prime\phi}T_{1}^{\prime},\qquad\tilde{T}_{2}^{\prime}=\tilde{T}_{1}^{\prime L_{1}}=T_{2}^{\prime}$ (35)
which satisfy $\mathcal{H}^{\prime}=\tilde{T}_{i}^{\prime}\mathcal{H}^{\prime}\tilde{T}_{i}^{\prime-1}$. Note that $L_{2}$-dependence in $\tilde{T}_{1}^{\prime}$ is again removed.
The adiabatic flux insertion operator is introduced as in the same way of Eq. (15) as
$F_{i}^{\prime}(\Phi_{i})=\mathcal{T}\exp\left[-{\rm i}\int_{0}^{T}\mathcal{H}_{i}^{\prime}(\Phi_{i}\tfrac{t}{T}){\rm d}t\right]$ (36)
and the Laughlin-type relation becomes
$F_{1}^{\prime}(\Phi)]^{-1}U_{2}^{\prime}F_{1}^{\prime}(\Phi)=U_{2}^{\prime}{\rm e}^{+{\rm i}\nu\Phi}.$ (37)
Since, in the system with SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect of the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation of Eq. (33), the magnetic translation operator corresponding to Eq. (17) is
$\bar{T}_{1}^{\prime}=F_{2}^{\prime}(2\pi L_{2}\phi)T_{1}^{\prime}=F_{1}^{\prime}(2\pi\phi)T_{1}^{\prime}.$ (38)
Hence, both Eqs. (34) and (38) become free of $L_{2}$ dependence. This is the essential advantage of SBCs. Therefore, we obtain the relation
$\bar{T}_{1}^{\prime}U_{2}^{\prime}\bar{T}_{1}^{\prime-1}=U_{2}^{\prime}{\rm e}^{{\rm i}2\pi(\phi\nu-\rho)},$ (39)
without any geometric parameters in contrast to Eq. (18). This means that the LSM-type argument under SBCs leads to the QHE condition $\phi\nu-\rho\in\mathbb{Z}$ directly, without passing through the intermediate relation $L_{2}(\phi\nu-\rho)\in\mathbb{Z}$.
–Summary and Discussion–
In this work, we have derived the condition for the appearance of the quantum Hall effect (QHE) in 2D lattice systems $\phi\nu-\rho\in\mathbb{Z}$, by means of a Lieb-Schultz-Mattis (LSM)-type and Laughlin-type argument. By introducing the spiral boundary conditions (SBCs), we have successfully eliminated the artificial dependence on the system width $L_{2}$ that arises in the conventional argument based on periodic boundary conditions (PBCs). Under the SBC, the 2D lattice can be mapped onto an extended 1D chain, so that explicit dependence on the cross section can be avoided, and the effects of flux insertion in different directions can be treated on equal footing. This construction provides a more transparent physical picture of the quantization of the Hall conductance in lattice fermion systems.
The key to this approach is that the spatial directions of the external force and the response can be manipulated by a factor of the system size as in Eqs. (33)-(35), and (38), thereby allowing the flux insertion process to be represented without introducing any redundant parameters. Consequently, the condition $\phi\nu-\rho\in\mathbb{Z}$ emerges directly from the magnetic translation symmetry, demonstrating that the quantization of the Hall conductance can be understood purely from symmetry and topology, without relying on explicit band-structure calculations.
Our approach unifies the LSM argument for interacting systems and Laughlin’s gauge argument for the QHE, and provides a useful framework for analyzing correlated topological phases in lattice models. Since the present formalism is formulated entirely in terms of the many-body wave function and the polarization operators, it can be straightforwardly extended to systems with interactions, disorder, or nontrivial lattice geometries. Applications to Chern insulators, fractional Chern insulators, and topological superconductors under SBCs will be interesting future problems.
–Acknowledgment–
The authors thank Y. Imamura and M. Oshikawa for discussions. M. N. acknowledges the research fellow position of the Institute of Industrial Science, The University of Tokyo. M. N. is supported partly by MEXT/JSPS KAKENHI Grant No. JP20K03769.

### v2

@@
 Abstract
-We derive the condition for the occurrence of the quantum Hall effect in two-dimensional lattice systems, expressed as $\phi\nu-\rho\in\mathbb{Z}$, where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively.
+We derive the condition for the occurrence of the integer quantum Hall effect in two-dimensional lattice systems with interactions, expressed as $\phi\nu-\rho\in\mathbb{Z}$, where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively.
 By employing spiral boundary conditions, which treat the system as an extended one-dimensional chain, this condition is obtained directly through a Lieb-Schultz-Mattis-type and Laughlin-type argument.
-This approach improves upon the preceding work based on conventional periodic boundary conditions, where the condition was derived indirectly containing redundant system-size dependence.
-The key to this approach is that the spatial directions of the external force and the response can be manipulated by a factor of the system size.
+This approach improves upon the preceding work based on conventional periodic boundary conditions, where the condition was derived indirectly with redundant system-size dependence.
+The key to this approach is that the spatial directions of the external force and the response can be systematically controlled by a factor of the system size.
 –Introduction–
@@
 The theory of polarization has also been applied to various systems[20, 21], and also to the calculation of topological numbers of 2D lattice systems by applying Laughlin’s arguments of flux insertion[22].
-Lu, Ran and Oshikawa derived the conditions for the appearance of the QHE in lattice systems with interactions by combining the LSM-type argument with Laughlin-type argument as[23, 24]
+Lu, Ran, and Oshikawa derived the condition for the appearance of the QHE in lattice systems with interactions by combining the LSM-type argument with Laughlin-type argument as[23, 24]
 $\phi\nu-\rho\in\mathbb{Z},$ (1)
@@
 In their theory based on periodic boundary conditions (PBCs), the condition for QHE appears indirectly passing through the intermediate relation which depends on an artificial geometric parameter.
-Therefore, improvement of this proof is desirable to eliminate this redundancy.
+Therefore, it is desirable to improve the proof to eliminate this redundancy.
 In this Letter, we reformulate the argument using spiral boundary conditions (SBCs)[16, 25, 26, 27], which allow a 2D system to be regarded as an extended 1D chain (see Fig.
@@
 This formalism provides a unified and more transparent understanding of the QHE from the standpoint of symmetry and topology.
+Figure 1: (Color) Boundary conditions for 2D lattices: (a) Conventional periodic boundary conditions (PBCs), (b) Spiral boundary conditions (SBCs), where the system is described as an extended 1D chain.
+Fluxes $\Phi_{i}$ for 2D lattices are applied as $\mathcal{H}_{i}(\Phi)=U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$.
+For SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect as the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation $U_{1}^{\prime}=U_{2}^{\prime L_{2}}$.
 In the rest of this Letter, we explain the argument by Ref.
@@
 Throughout this Letter, the electron charge, the reduced Planck constant, and the lattice constant are set to be unity as $e=\hbar=a=1$.
-Figure 1: Several boundary conditions for 2D square lattices with $L_{1}\times L_{2}$ sites: (a) Conventional periodic boundary conditions (PBCs), (b) Spiral boundary conditions (SBCs).
-For the systems with SBCs, the lattices are labeled as extended 1D chains.
-$\bm{A}_{i}$ and $\bm{A}_{i}^{\prime}$ are primitive vectors.
 –Conventional periodic boundary conditions–
@@
 $\mathcal{H}=-t\sum_{m,n}(c^{\dagger}_{m+1,n}{\rm e}^{{\rm i}\theta_{m,n}^{1}}c^{\mathstrut}_{m,n}+c^{\dagger}_{m,n+1}{\rm e}^{{\rm i}\theta_{m,n}^{2}}c^{\mathstrut}_{m,n}+\mbox{H.c.})+V(\{n_{m,n}\}),$ (2)
-where $c^{\mathstrut}_{m,n}$ ($c_{m,n}^{{\dagger}}$) is the annihilation (creation) operator at $(m,n)$ site, $t$ is the hopping amplitude, $V(\{n_{m,n}\})$ is the fermion interactions consists of the number operator $n_{m,n}\equiv c_{m,n}^{{\dagger}}c_{m,n}^{\mathstrut}$.
+where $c^{\mathstrut}_{m,n}$ ($c_{m,n}^{{\dagger}}$) is the annihilation (creation) operator at $(m,n)$ site, $t$ is the hopping amplitude, $V(\{n_{m,n}\})$ represents the fermion interaction and is written in terms of the number operator $n_{m,n}\equiv c_{m,n}^{{\dagger}}c_{m,n}^{\mathstrut}$.
 The gauge field is introduced in the Landau gauge as
 $\theta^{1}_{m,n}=0,\qquad\theta^{2}_{m,n}=2\pi\phi m,$ (3)
-which give the magnetic flux $B=-\frac{2\pi\hbar}{ea^{2}}\phi$ with
+which gives the magnetic flux $B=-\frac{2\pi\hbar}{ea^{2}}\phi$ with
 $\phi=\frac{1}{2\pi}\sum_{\Box}\theta_{m,n}=\frac{p}{q},$ (4)
@@
 $\sum_{\Box}\theta_{m,n}=\theta_{m,n}^{1}+\theta_{m+1,n}^{2}-\theta_{m,n+1}^{1}-\theta_{m,n}^{2}.$ (5)
-Figure 2: The gauge field $\theta^{\alpha}$ in a square lattice in a magnetic field with (a) periodic boundary conditions, and (b) spiral boundary conditions.
 In the absence of a magnetic field ($\phi=0$), the Hamiltonian possesses translational symmetry by one lattice spacing along both directions, $\mathcal{H}=T_{i}^{\mathstrut}\mathcal{H}T_{i}^{-1}$.
@@
 Therefore, $L_{1}$ is selected so that $L_{1}/q\in\mathbb{Z}$.
-On the other hand, a similar symmetry is recovered as $\mathcal{H}=\tilde{T}_{i}^{\mathstrut}\mathcal{H}\tilde{T}_{i}^{-1}$ by introducing the magnetic translation operators
+On the other hand, a similar symmetry is satisfied as $\mathcal{H}=\tilde{T}_{i}^{\mathstrut}\mathcal{H}\tilde{T}_{i}^{-1}$ by introducing the magnetic translation operators
 $\tilde{T}_{1}=U_{2}^{L_{2}\phi}T_{1},\qquad\tilde{T}_{2}=T_{2}.$ (6)
@@
 where $\hat{\rho}$ is the density operator, and $L_{2}$ represents the cross section of the 2D system. As in the LSM argument for $d\geq 2$, $L_{2}$ and $q$ should be chosen to be coprime.
-Figure 3: Sublattice structure of the system $(L_{1},L_{2},q)=(6,4,3)$ with a magnetic field ($\phi=p/q$). Here, $L_{1}/q$ is chosen to be an integer.
+Figure 2: (Color) The gauge field $\theta^{\alpha}$ in a square lattice in a magnetic field with (a) periodic boundary conditions, and (b) spiral boundary conditions.
+Figure 3: (Color) Sublattice structure of the system $(L_{1},L_{2},q)=(6,4,3)$ with a magnetic field ($\phi=p/q$). Here, $L_{1}/q$ is chosen to be an integer.
 Now we consider the quantum Hall effect in this lattice system. The quantized Hall conductivity is characterized by the Chern number $\nu$ and the density $\rho$ as
@@
 $\nu=-\frac{1}{\Phi}\mathrm{Im\,ln\,}z(\Phi).$ (13)
+This is also an expression of Laughlin’s argument for QHE [4] in lattice systems.
+It states that inserting a unit flux $\Phi=2\pi h/e$ in the $x_{2}$ direction pumps an integer number of fermions in the $x_{1}$ direction when $\sigma_{xy}$ is quantized.
 The adiabatic flux insertion is expressed as
@@
 $F_{i}(\Phi_{i})=\mathcal{T}\exp\left[-{\rm i}\int_{0}^{T}\mathcal{H}_{i}(\Phi_{i}\tfrac{t}{T}){\rm d}t\right],$ (15)
-where $\mathcal{T}$ denotes time ordering, $T$ is a sufficiently long period, and $\mathcal{H}_{i}(\Phi)\equiv U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$.
+where $\mathcal{T}$ denotes time ordering, $T$ is a sufficiently long period compared with the inverse of the energy gap, and
+$\mathcal{H}_{i}(\Phi)\equiv U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}.$ (16)
 Equation (12) then yields the operator relation
 $F_{2}(\Phi)]^{-1}U_{1}F_{2}(\Phi)=U_{1}{\rm e}^{-{\rm i}\nu\Phi}.$ (17)
-This is also an expression of the Laughlin’s argument for QHE. Since $F_{i}(2\pi)$ plays a similar role to $U_{i}$, we can introduce a magnetic translation operator as an analogy of Eq. (6),
+Since $F_{i}(2\pi)$ plays a similar role to $U_{i}$, we can introduce a magnetic translation operator as an analogy of Eq. (6),
 $\bar{T}_{1}=F_{2}(2\pi L_{2}\phi)T_{1},$ (18)
@@
 $\bar{T}_{1}U_{1}\bar{T}_{1}^{-1}=U_{1}{\rm e}^{{\rm i}2\pi L_{2}(\phi\nu-\rho)}.$ (19)
-Thus, following an LSM-type argument, the condition for $z(\Phi)\neq 0$ becomes
+Thus, following an LSM-type argument for the symmetry operation, the condition for $z(\Phi)\neq 0$ becomes
 $z(\Phi)$ $=\braket{\Psi_{2}(\Phi_{2})|\bar{T}_{1}U_{1}\bar{T}_{1}^{-1}|\Psi_{2}(\Phi_{2})}$ (20)
@@
 However, the appearance of the artificial parameter $L_{2}$ is undesirable, because $L_{2}$ must be chosen to be relatively prime to $q$ so that a redundancy remains.
-Figure 4: Fluxes $\Phi_{i}$ for 2D lattice applied as $\mathcal{H}_{i}(\Phi)=U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$.
-In the system with SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect of the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation $U_{1}^{\prime}=U_{2}^{\prime L_{2}}$.
 –Spiral boundary conditions–
@@
 The primitive vectors in the real space (see Fig.
-1(a)) and the reciprocal lattice space of 2D systems as
+4(a)) and the reciprocal lattice space of 2D systems as
 $\bm{A}_{1}$ $=(L_{1},0),$ $\qquad\bm{B}_{1}$ $=2\pi(\tfrac{1}{L_{1}},0),$ (25a)
@@
 $\frac{2\pi jn_{j}}{L_{1}}=\frac{2\pi x_{1}n_{\bm{x}}}{L_{1}}\quad(\mathrm{mod}\,2\pi).$ (30)
-For $i=2$, the exponent part of the polarization operator is confirmed as that of SBCs as
+For $i=2$, the exponent part of the polarization operator can be written in the SBC representation as
 $\sum_{\bm{x}}\bm{B}^{\prime}_{2}\cdot\bm{x}n_{\bm{x}}$ $=\sum_{m=1}^{L_{1}}\sum_{n=0}^{L_{2}-1}2\pi\frac{m+nL_{1}}{L_{1}L_{2}}n_{m,n}$ (31)
@@
 $U_{1}^{\prime}=\exp\left[{\rm i}\sum_{j=1}^{L}\frac{2\pi jn_{j}}{L_{1}}\right],\quad U_{2}^{\prime}=\exp\left[{\rm i}\sum_{j=1}^{L}\frac{2\pi jn_{j}}{L}\right].$ (33)
+Figure 4: (Color) Boundary conditions for 2D square lattices with $L_{1}\times L_{2}$ sites: (a) Periodic boundary conditions (PBCs), (b) Spiral boundary conditions (SBCs).
+For the systems with SBCs, the lattices are labeled as extended 1D chains.
+$\bm{A}_{i}$ and $\bm{A}_{i}^{\prime}$ are primitive vectors.
+When we cut the system (by green lines), the difference in the number of removed bonds in both boundary conditions is only one.
 In SBCs, the translation and polarization operators for different directions are related as
@@
 Then, the counterpart of Eq.(8) becomes
-$T_{1}^{\prime}U_{2}^{\prime}T_{1}^{\prime-1}=T_{1}^{\prime}U_{1}^{\prime 1/L_{2}}T_{1}^{\prime-1}={\rm e}^{-{\rm i}2\pi\hat{\rho}}U_{2}^{\prime}.$ (34)
+$T_{1}^{\prime}U_{2}^{\prime}T_{1}^{\prime-1}={\rm e}^{-{\rm i}2\pi\hat{\rho}}U_{2}^{\prime}.$ (35)
 We choose $U_{2}^{\prime}$ rather than $U_{1}^{\prime}$ here because it eliminates the $L_{2}$-dependence from this relation.
-This was introduced for an extention of the LSM theorem to higher dimensions[16].
+This was introduced for an extension of the LSM theorem to higher dimensions[16].
 The magnetic translation operators are
@@
 $F_{i}^{\prime}(\Phi_{i})=\mathcal{T}\exp\left[-{\rm i}\int_{0}^{T}\mathcal{H}_{i}^{\prime}(\Phi_{i}\tfrac{t}{T}){\rm d}t\right]$ (37)
-and the Laughlin-type relation becomes
+and the Laughlin-type relation (17) becomes
 $F_{1}^{\prime}(\Phi)]^{-1}U_{2}^{\prime}F_{1}^{\prime}(\Phi)=U_{2}^{\prime}{\rm e}^{+{\rm i}\nu\Phi}.$ (38)
-Since, in the system with SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect of the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation of Eq.
-(33), the magnetic translation operator corresponding to Eq.
+In Eq.
+(37), the Hamiltonian with a flux $\mathcal{H}_{i}^{\prime}(\Phi_{i})$ is described similarly to the case of PBC, however, the definition Eq.
+(16) for SBCs means that effects of the fluxes $\Phi_{1}$ and $\Phi_{2}$ are no longer independent (see also Fig.
+1(b)), because $U_{1}^{\prime}$ and $U_{2}^{\prime}$ have the same phase variable except for their coefficients as in Eq.
+(33).
+Therefore, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect as the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation of Eq.
+(34), so that the magnetic translation operator corresponding to Eq.
 (18) is
@@
 –Summary and Discussion–
-In this work, we have derived the condition for the appearance of the quantum Hall effect (QHE) in 2D lattice systems $\phi\nu-\rho\in\mathbb{Z}$, by means of a Lieb-Schultz-Mattis (LSM)-type and Laughlin-type argument.
+In this work, we have derived the condition for the appearance of the integer quantum Hall effect (QHE) in 2D lattice systems with arbitrary interactions, namely $\phi\nu-\rho\in\mathbb{Z}$, by means of a Lieb-Schultz-Mattis (LSM)-type and Laughlin-type argument.
 By introducing the spiral boundary conditions (SBCs), we have successfully eliminated the artificial dependence on the system width $L_{2}$ that arises in the conventional argument based on periodic boundary conditions (PBCs).
-Under the SBC, the 2D lattice can be mapped onto an extended 1D chain, so that explicit dependence on the cross section can be avoided, and the effects of flux insertion in different directions can be treated on equal footing.
+Under the SBC, the 2D lattice can be mapped onto an extended 1D chain, so that explicit dependence on the system’s cross section is avoided, and the effects of flux insertion in different directions are treated on equal footing.
 This construction provides a more transparent physical picture of the quantization of the Hall conductance in lattice fermion systems.
-The key to this approach is that the spatial directions of the external force and the response can be manipulated by a factor of the system size as in Eqs.
+The key to this approach is that the spatial directions of the external force and the response can be controlled by a factor of the system size as in Eqs.
 (34)-(36), and (39), thereby allowing the flux insertion process to be represented without introducing any redundant parameters.
 Consequently, the condition $\phi\nu-\rho\in\mathbb{Z}$ emerges directly from the magnetic translation symmetry, demonstrating that the quantization of the Hall conductance can be understood purely from symmetry and topology, without relying on explicit band-structure calculations.
+Although we emphasize the advantages of the 1D description by SBCs, the system remains physically equivalent to the 2D lattice, and hoppings in the $x_{2}$ direction still appear as long-range hopping terms $c^{\dagger}_{j+L_{1}}{\rm e}^{{\rm i}\theta_{j}^{2}}c^{\mathstrut}_{j}+\mbox{H.c.}$ in the 1D chain representation.
+Thus, cutting the torus under SBCs removes not only the bond on the 1D chain, but also $L_{1}$ long-range bonds as shown in Fig.
+4.
+Therefore, compared with PBCs, the number of removed bonds differs only by one, which does not affect the physics in the thermodynamic limit.
+Hence, the edge-state properties under the two boundary conditions are essentially the same, and Laughlin’s relation holds in both cases.
 Our approach unifies the LSM argument for interacting systems and Laughlin’s gauge argument for the QHE, and provides a useful framework for analyzing correlated topological phases in lattice models.
-Since the present formalism is formulated entirely in terms of the many-body wave function and the polarization operators, it can be straightforwardly extended to systems with interactions, disorder, or nontrivial lattice geometries.
-Applications to Chern insulators, fractional Chern insulators, and topological superconductors under SBCs will be interesting future problems.
+Since the present formalism is formulated entirely in terms of the many-body wave function and the polarization operators, it can be straightforwardly extended to systems with disorder or nontrivial lattice geometries, and also to the fractional QHE[23].
+In addition to the present method for the QHE condition, an alternative formulation has also been proposed by Matsugatani et al.[24] in terms of many-body Chern number based on Berry connections [1, 29] and point-group symmetries.
+It would be also interesting to reformulate their framework using SBCs.
 –Acknowledgment–